if the perimeter of a rectangle is 150 find the area
Answer by mananth(16946)

Perimeter= 150
square edge be x
all are equal
so 4x=150
x=37.5 the side length
..
Area = L*W
But it is a square
so Area = L*L
Are = 37.5*37.5
Area = 1406.25